Prove that 2n>4n for n≥5.



1
Expert's answer
2021-06-07T13:34:37-0400

Prove for "n=5" :

"2^5=32>20=4\cdot5"

We prove by induction, let be true for "n-1"

"2^n=2\cdot2^{n-1}\stackrel{\text{True for n-1}}{>}2\cdot4(n-1)=4n+(4n-8)"

Because "4n-8>0" for "n\geq5" :

"4n+(4n-8)>4n"

So, we prove that "2^n>4n" for "n\geq5"